OpenStudy (phebe):

HELP!!! Straight answers ok Which radical expression is in simplified form? im going to post the answer choices too

OpenStudy (a1234):

No straight answers.

OpenStudy (anonymous):

we need the questions

Directrix (directrix):

OpenStudy values the Learning process - not the ‘Give you an answer’ process •Don’t post only answers - guide the asker to a solution. •Have homework problems? We can help you - as long as you are looking to learn how to get your solution (and not just the answer). http://openstudy.com/code-of-conduct @Phebe

OpenStudy (anonymous):

...

OpenStudy (phebe):

A-\[\frac{ 11y }{ \sqrt{3}? }\] B-\[\frac{ \sqrt{6} }{ 5y? }\] C-\[\frac{ \sqrt{17} }{ ?4 }\] D-\[\frac{ \sqrt{25} }{81 ? }\]

OpenStudy (phebe):

the question is at the top

OpenStudy (phebe):

foreget that question mark on C that was an accident k

OpenStudy (anonymous):

for the second one is it supposed to be (√6)/(5y) ?

OpenStudy (phebe):

yes

OpenStudy (anonymous):

a

OpenStudy (a1234):

You can eliminate A since a radical is not supposed to be in the denominator. D has a perfect square under the radical. Try deciding between B and C @Phebe

OpenStudy (phebe):

thanx @Warriorslayer u sure

OpenStudy (phebe):

ok @a1234 are u sure also

OpenStudy (a1234):

@Warriorslayer is wrong.

OpenStudy (anonymous):

if not then i would say c

OpenStudy (phebe):

ok thanx @a1234

OpenStudy (phebe):

ok i'll pick C

OpenStudy (anonymous):

the answer is c

OpenStudy (anonymous):

beacuse if you get rid of the mistake you made its simplified

OpenStudy (anonymous):

im not the best at geometry im great at history though

OpenStudy (anonymous):

but thats just my answer you might want someone to second that

OpenStudy (phebe):

thanx EVEERYONE

OpenStudy (phebe):

it wasn't C it was B

OpenStudy (a1234):

I can't see why B is wrong...C is also right.

OpenStudy (phebe):

B is the right answer

OpenStudy (anonymous):

im soooooooooosorry. you never answered my question i asked earlier so i guessed that that menat no. :( sorry again.

OpenStudy (a1234):

If 4 was under the radical in C, then it would be wrong.

OpenStudy (phebe):

its ok

OpenStudy (a1234):

Was it under the radical @Phebe? Just checking.

OpenStudy (phebe):

yes

OpenStudy (anonymous):

she closed the question

OpenStudy (a1234):

Well, sqrt4 is 2, so you can see why C is wrong :).

OpenStudy (anonymous):

wait 4 was under a radical?!

OpenStudy (anonymous):

omg! well then duh it would be b! :(

OpenStudy (anonymous):

If you would of put the radical form over the four then I wouldve said c. I got a 100% on my radical, and square root test. LOVE math

OpenStudy (a1234):

You mean B, right @MeganChase?

OpenStudy (anonymous):

Yeah...lol